OpenStudy (anonymous):

Haven't done this since 6th grade and I'm a freshman in college, find the equation of a line that passes through (2,12) and is parallel to the graph of y=2x+4, can someone explain please? Thanks!

OpenStudy (triciaal):

with practice it will come back to you. parallel lines have the same slope the equation of a line is such that y is a function of x y is dependent on the value of x and the slope is the rate of change. this is the ratio of how much y changes for the corresponding change in x y = m x + b m is the slope of the line and b is the value of y when the value of x = 0 b is also known as the y-intercept what is the slope of the given line?

OpenStudy (triciaal):

y = m x + b y = 2 x + 4 what is m ?

OpenStudy (anonymous):

m=2

OpenStudy (triciaal):

good

OpenStudy (triciaal):

we know that the second line passes through the point (2, 12) this is the (x, y) the line is parallel to the first line so m is the same = 2 and y = m x + b so we can put the numbers in and find b what is b?

OpenStudy (anonymous):

12?

OpenStudy (triciaal):

no (2, 12) means x = 2 and y = 12 a point is always (x, y) y = m x + b you have y, m and x replace the letters with the numbers and you only have one unknown in the equation to solve that's b

OpenStudy (anonymous):

so would it be y=-2x+12

OpenStudy (triciaal):

y = m x + b 12 = 2(2) + b

OpenStudy (anonymous):

b=8

OpenStudy (triciaal):

yes so for the parallel line to the one given the slope m = 2 and b = 8 and to find any point the point is (x, y) the equation is written y = m x + b replace m and b with the numbers

OpenStudy (anonymous):

wow, thank you that makes a lot more sense. I really appreicatr your help.

OpenStudy (triciaal):

no problem

OpenStudy (triciaal):